OpenStudy (anonymous):

A and X is 4x4 matrix. AX=E-2X. Find X.

OpenStudy (anonymous):

Add 2X both the sides first...

OpenStudy (anonymous):

\[AX + 2X = E - 2X + 2X\] \[AX + 2X = E\]

OpenStudy (anonymous):

Can you tell me what is E here?? A matrix or something else..??

OpenStudy (anonymous):

When you will take common X from AX + 2X then you will be using Identity matrix that is I: \[X(A + 2I) = E\]

OpenStudy (anonymous):

im not sure the answer is X=(A+2E)inverse(-1)

OpenStudy (anonymous):

want to know why A+2E

OpenStudy (anonymous):

And now to find X, you will take inverse of \((A + 2I)\) So, you will get : \[X = (A+2I)^{-1} \cdot E\]

OpenStudy (anonymous):

Do you know we can add two matrices having same rows and columns ??

OpenStudy (anonymous):

Suppose you have A and B matrices of 3x3 dimension: You can add, subtract these matrices.. Right ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Remember one thing : You can add or subtract two matrices if they have same dimensions.. But you cannot add a matrix with a number like 2 ; I show you : A + 2 = ???

OpenStudy (anonymous):

Here, you can't add a matrix with a number: However, you can add two matrices..

OpenStudy (anonymous):

To make 2 as matrix, we multiply it with identity matrix (E or I) having same dimension of the matrix we are adding it to...

OpenStudy (anonymous):

Getting ??

OpenStudy (anonymous):

so is E the identity matrix?

OpenStudy (anonymous):

@hartnn will help you here... Electricity has gone here... Ha ha ha... Here, In India, Identity matrix is shown by I...

OpenStudy (anonymous):

i study in China. Math here is wierd and hard

OpenStudy (anonymous):

i kind of get what you said

OpenStudy (anonymous):

AX=E-2X to get matrix 2 that makes (A+2E)X=E then X=(A+2E)inverse

hartnn (hartnn):

u got that correct @mxgforce